The active element of a certain laser is made Glas rod 30cm long And 1.50cm in diameter. What will be the increase in it volume if the temperature of the rod is increase 70.0degree
1
Expert's answer
2020-10-30T13:19:22-0400

The increase in it volume is given by the following expression:


"\Delta V = \alpha_V\cdot V\cdot \Delta T"

where "\alpha_V = 25.5\times 10^{-6}K^{-1}" is the volumetric thermal expansion coefficient of glass, "V" is the initial volume and "\Delta T = 70\degree C = 70K" is the  increase of temperature. The initial volume of the rode is:


"V = \dfrac{\pi d^2L}{4}"

which is the volume of a cylinder with a diameter "d = 1.5cm = 0.015m" and length "L = 30cm = 0.3m". Substituting the expression for the volume into the first equation, obtain:


"\Delta V = \alpha_V\cdot \dfrac{\pi d^2L}{4}\cdot \Delta T \\
\Delta V = 25.5\times 10^{-6}\cdot \dfrac{\pi\cdot 0.30\cdot 015^2\cdot }{4}\cdot 70 \approx 9.46\times 10^{-8}m^3"

Answer. "9.46\times 10^{-8}m^3".